June 9, 2010 Tracking CDC/FDC
Meeting Time and Place
Wednesday June 9, 2010 at 11:00am At Jefferson Lab, the meeting will be held in room F326-327
To connect from the outside, please use ESNET
1.) ESNET: 8542553
+1-800-377-8846 : US
+1-888-276-7715 : Canada
+1-302-709-8424 : International
then enter participant code: 77438230# (remember the "#")
- FDC Update logbook
- CDC Update
Present at the meeting: C. Meyer, Y.Van Haarlem and N. Jarvis from CMU, S.Taylor, D. Lawrence, B. Zihlmann, T. Whitlatch and L. Pentchev from Jefferson Lab.
- There were no announcements.
- Lubomer reported on the status of the FDC.
- The UVA stringing equipment is now at JLab and they are looking into setting it up in F116. UVA will continue to work on the position and tension measurements and have kept that part of the equipment at UVA.
- Cosmic tests continue using the new preamplifier and the new FADC (125). There is an apparent high-rate issue with the FADC-125 which is being looked at. Using all the information available from the FADC, he is able to achieve 100-micron resolution with the cathodes. While this is probably better than is needed with the multiple scattering, it certainly would help pattern recognition.
- Work with GARFIELD is going on to try to optimize the field settings.
- Beni reported on his work.
- The source of common noise that was seen on the system appears to been coming from the UPS that was driving the CAEN HV unit. This has been separated from the signal chain and the noise is gone.
- He showed a 97% efficiency for the wires.
- Tim noted that there are four applicants for the FDC work coordinator and the the PR for the clean room had gone out.
- Curtis reported on the CDC.
- We went through the Status Report that was sent to Jefferson Lab last friday on where things stand with construction.
- He noted that the Medelec crimp pins arrived safely at CMU. They appear clean and are much easier to string than the older CLAS pins were.
- Tim noted that the copper ones in the sample were from acumetrics. It is not clear if we will have those plated or not.
- Curtis then pointed people to a CMU site where photos of the work at CMU are available.
- Naomi reported on the ASIC
- There were a number of issues in getting set up with the new preamp, and she believes that the saturation was distorting her perception of what the scope traces looked like.
- She set up both cosmics and iron-55 into the data acquisition and we are not sure what is going on with the preamplifier. With the chamber using the same gas and the same high-voltage settings, runs were made with both the old preamplifier and the new preamplifier. The old has a maximum output of 1.0V, while in the new one, it appears to have been reduced to 0.5V. This was unexpected!
- We also note that the new preamplifier is supposed to have 1/3 of the gain of the older one. This is to allow for a larger dynamic range of signals and is important for both position resolution and dE/dx.
- We are observing that the gain of the new preamplifier is a factor of 3 LARGER than the old one, not smaller! Thus, the dynamic range has gone down by a factor of six, and not up by a factor of three. If true, this is very bad.
- A meeting has been set up on Monday at 10:00am to discuss this with Fernando.
- A follow-up discussion came up in trying to get information on the CMU wire-tension measurements. We post here a few links that might be useful.
- Reinhard Schumacher wrote an article on the CMU method Roth & Schumacher NIM 369(1996)215-221,(DOI Link).
- The instructions for how to measure the wire tension at CMU.
- We also recently got some information from Mac Mestayer on a new method to measure wire tensions that they will be using for CLAS12.File:Tension-1.pdf as well as a File:James Haus Senior Thesis-1.pdf describes this in detail.